OpenStudy (anonymous):

A cell phone company offers two different monthly text-messaging plans as shown in the table below. For what number of text messages will both plans cost the same? HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Messaging Plan Monthly Fee Fee per Message Plan A $35.00 $0.10 Plan B $18.00 $0.30

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

Can you make equations for the total cost of each plan?

OpenStudy (anonymous):

Ok, for Plan A the total cost can be represented by\[C_{a}=$35+$0.10 \times T\]where T is the number of texts. Can you make an equation for Plan B?

OpenStudy (anonymous):

u let me see

OpenStudy (anonymous):

what does Ca stand for does a stand for plan A

OpenStudy (anonymous):

The cost of Plan A, yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Cb=$18+ $0.30*T

OpenStudy (anonymous):

Yes! So there's going to be a value of T that makes Ca=Cb. So just solve\[18+0.30 \times T=35+0.10 \times T\]and you're done!

OpenStudy (anonymous):

ALRIGHT SO I GOT T=85

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thank you you always break them down and make the ones that were hard super easy

OpenStudy (anonymous):

You're very welcome. :)